A gas absorbs a photon of 355 nm and emits at two wavelengths.If one of the emissions is at 680 nm, the other is at:-
a.325 nm    b.743nm

Grade:

1 Answers

Mohd abdul rafey
4 years ago
1/∂=1/∂1+1/∂2
so
1/355=1/680+1/∂2
1/∂2=325/680*355
answer is 743nm
hope you find this solution useful